[Brahms-l] Shift report 20050225 16:00-24:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Fri Feb 25 2005 - 23:53:46 EST
Shifter leader was HHD
Shifters were 


1. MCR dumped the first beam before time to steer for STAR. This meant it took some time before the next beam was up and running. No technical issues.

2. Data runs taken with FS@2.3B1723:
   13484(104k), 13487(244k), 13488(436k), 13489(283k)
   Non-TPC pedestal run: 13485
   TPC pedestal run: 13486

3. No problems (!).

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Fri Feb 25 23:54:09 2005

This archive was generated by hypermail 2.1.8 : Fri Feb 25 2005 - 23:54:25 EST